A sample of sulphur dioxide occupies a volume of 652mL at 40C and 720mmHg. What volume will the sulphur dioxide occupy at STP?


1
Expert's answer
2022-02-27T19:45:02-0500

Use the ideal gas law to find the moles of sulfur dioxide. Then use the fact that at STP, 1 mole of gas occupied 22.4 liters.


PV = nRT

P = 720 mm Hg x 1 atm / 760 mm Hg = 0.947 atm

V = 652 ml = 0.652 L

n = ?

R = 0.0821 Latm/Kmol

T = 40º + 273 = 313K

Solving for n, we have...

n = PV/RT = (0.947)(0.652) / (0.0821)(313)

n = 0.0240 moles SO2

At STP..

0.0240 mols x 22.4 L/mol = 0.538 L

Volume of sulphur dioxide at STP is;

= 0.538L